OpenStudy (anonymous):

Marrie is logging the number of radioactive isotope she observes every minute. She has determined her function to be f(x) =-4x +120. Based on the situatuion above describe and neccesary restrictions to the domain and range

OpenStudy (anonymous):

i guess the domain is minutes, so you would have to restrict it to positive numbers, as, (alas) you cannot go backwards in time

OpenStudy (anonymous):

also presumably you cannot have a negative number of isotopes, so the range has to be greater then zero this means also for the domain you have to solve \[-4x+120\geq 0\]

OpenStudy (anonymous):

*greater than or equal to zero not "greater then zero"

OpenStudy (anonymous):

Ok ..uh I dont understand anything lol ..can u break it down for me?

OpenStudy (anonymous):

do you know what "domain" means?

OpenStudy (anonymous):

the x and y values yess

OpenStudy (anonymous):

no that is not what it means the "domain" means the possible inputs of the function

OpenStudy (anonymous):

Oh :/ eh

OpenStudy (anonymous):

@satellite73

OpenStudy (agent0smith):

x is time. Time can not be negative, which is why x ≥ 0. Understand?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

it says to justify your anwer for the restictions

OpenStudy (agent0smith):

That's part of the domain, x ≥ 0. Remember DOMAIN is x values, not y values. The other part of the domain, as satellite said, you can't have a negative number of isotopes. So for f(x) =-4x +120, the \[\large -4x+120\geq 0 \] which you need to solve.

OpenStudy (anonymous):

x = 30

OpenStudy (agent0smith):

Not quite correct. There isn't an equals sign here \[\large -4x+120\geq 0\] So there won't be an equals sign in your answer.

OpenStudy (anonymous):

x>30 ..sorry

OpenStudy (agent0smith):

Not quite but closer. \[\large -4x+120\geq 0\] first subtract 120 from both sides \\[\large -4x \ge -120\] now divide both sides by -4. What happens when you divide both sides of an inequality by a negative number?

OpenStudy (anonymous):

becomes poitive

OpenStudy (agent0smith):

What happens to the sign of the inequality when you divide both sides by a negative?

OpenStudy (anonymous):

it flips?

OpenStudy (anonymous):

x<30

OpenStudy (agent0smith):

Correct. So we had x ≥ 0 and now we have x =< 30. Which we can combine, since x is greater than or equal to 0, and less than or equal to 30 \[\Large 0 \le x \le 30\] That's our domain.